OpenStudy (anonymous):

A standard piece of printer paper has the dimensions 8.5" by 11". What is the ratio of the longer side to the shorter to the nearest hundredth? A. 1.61 B. 1.29 C. 1.15 D. 0.77

OpenStudy (anonymous):

L/S = 11/8.5 = 1.29

OpenStudy (anonymous):

thanks, glad you're back

OpenStudy (anonymous):

Temporarily!

OpenStudy (anonymous):

and uw!